OpenStudy (jhannybean):

A proton is released from rest in a uniform electric field of magnitude \(1.08 \times 10^5 \frac{N}{C}\) . Find the speed of the proton after it has traveled \(a. ~ 1.00 ~ cm~,~ b. ~ 10.0~cm\)

OpenStudy (anonymous):

Ummm, energy methods?

OpenStudy (perl):

you can use the equation F = q*E

OpenStudy (jhannybean):

I know that in finding the speed of a proton, given that it starts at \(V_i = 0 ~m/s\) would be \[v_f^2 = v_i^2 +2a\Delta x\]

OpenStudy (anonymous):

How many C is a proton? Multiply that to get N, which is a measure of force. Multiply that by each distance to get the work. Then use kinetic energy. \[ Fd = \frac 12 mv^2 \]

OpenStudy (perl):

substitute that for force

OpenStudy (jhannybean):

Oh, because the electric field is \(\vec E =\dfrac{\vec F}{q}\)?

OpenStudy (perl):

yes

OpenStudy (anonymous):

I was just using dimensional analysis to guess that \(q\mathbf E =\mathbf F\).

OpenStudy (jhannybean):

Ohh..

OpenStudy (anonymous):

Because \(\frac NC \cdot C = N\) and we wanted \(N\).

OpenStudy (jhannybean):

Ohh, I wasn't familiar with units.

OpenStudy (jhannybean):

Now it makes a little more sense.

OpenStudy (jhannybean):

then... \[\sum F = m\vec a = \vec F q \] \[\vec a = \frac{\vec Fq}{m}\]

OpenStudy (jhannybean):

And i'm guessing the units would cancel out somehow giving us... m/s^2?

OpenStudy (jhannybean):

Oh, mistake.

OpenStudy (jhannybean):

\[\vec a = \frac{\vec E q}{m} = \frac{\frac{N}{C} \cdot C}{kg} = \frac{\dfrac{kg\cdot m}{s^2}}{kg}=\frac{m}{s^2}\]

OpenStudy (jhannybean):

Hmm..

OpenStudy (jhannybean):

\[ v_f^2 = v_i^2 +2 a \Delta x~,~ \Delta x = 0.0100 ~m\]\[v_f =\sqrt{v_i^2 +2\left(\frac{\vec Eq}{m}\right)(0.0100~m)}\]

OpenStudy (jhannybean):

Right? @wio

OpenStudy (anonymous):

Yeah

OpenStudy (jhannybean):

Alright, and... um.. \(\vec E\) is given. But what is q?...

OpenStudy (anonymous):

The charge of a proton...

OpenStudy (jhannybean):

Oh, haha the electrical charge.... ~,~ \(1.08 \times 10^{-19} C\)... Duh.

OpenStudy (jhannybean):

Thanks!

OpenStudy (anonymous):

Hmm, yeah. Be glad the field was uniform

OpenStudy (jhannybean):

What if it wasn't? What would you do then?

OpenStudy (jhannybean):

The same thing but calculate each part? My guess.

OpenStudy (anonymous):

Then you'd have to know what you're doing, lol

OpenStudy (anonymous):

If field varied with respect to your distance from it, like a point charge, then you'd have to integrate to find the work done.

OpenStudy (jhannybean):

I see.